Problem 3 If logo If Togo3=0.5 and 109.5=0.7, Findi al 10g. 33 b) 1096 43
Answer #1

Given:

log_b \;3=0.5 \; and \; log_b \; 5=0.7

a)

log_b \;\frac{3}{5}=log_b \;3-log_b \;5 .............. \because log_x\left ( \frac{m}{n} \right )=log_x\:m-log_x\:n

  =(0.5)-(0.7)

  =0.2

b)

log_b\:^4\sqrt 3=log_b\:3^{\frac{1}{4}}

  ={\frac{1}{4}}\: log_b\:3 ...........................   \because log_x\:m^n=n\:(log_x\:m)

  ={\frac{1}{4}}\: (0.5)

  ={\frac{1}{8}}

Note: If you want further assistance regarding this question feel free to ask in comment box.

.

PLEASE THUMBS'UP IF IT IS HELPFUL